Find the future value when investing $4,000 at the annual interest rate of 3.5% for 7 years compounded quarterly.
Answer by stanbon(75887)

Find the future value when investing $4,000 at the annual interest rate of 3.5% for 7 years compounded quarterly.
-------------
A(t) = P(1+(r/n))^(nt)
A(7) = 4000(1+(0.035/4))^(4*7)
-------
A(7) = 4000*1.2763 = $5105.04
================================
Cheers,
Stan H.
===================